OpenStudy (javsteratl):

Find an equation for the line that passes through (-4, 8) and (3, -7). What is the slope? Where does the line intersect the x-axis and the y-axis?

OpenStudy (mathmale):

There is a slope formula that you really do need to learn by heart. Look up "slope" on the 'Net or in your textbook. Once you have the slope, use the "slope-intercept form of the equation of a straight line to complete this problem. This equation has the form y=mx+b. substitute your slope for "m" in this formula. Now choose one of the two given points. It has an x-coordinate and a y-coordinate. Substitute your y-coordinate for y in y=mx+b, and subst. your x-coord. for x in y=mx+b. Show your work. When you've reached this point, you'll be able to calculate the y-intercept, b.